# Density-based User Representation using Gaussian Process Regression for Multi-interest Personalized Retrieval

### v4

@@ 7 Conclusion and Discussion
In this paper, we introduced a density-based user representation model, GPR4DUR, marking the first application of Gaussian process regression for user modeling in multi-interest retrieval. This innovative approach inherently captures dynamic user interests, provides uncertainty-awareness, and proves to be more dimension-efficient than traditional point-based methods. We also establish a new evaluation protocol, developing new metrics specifically tailored to multi-interest retrieval tasks, filling a gap in the current evaluation landscape. Offline experiments validate the adaptability and efficiency of GPR4DUR, and demonstrate significant benefits relative to existing state-of-the-art models. Online simulations further highlight GPR4DUR’s ability to drive user interest exploration by recommendation algorithms that can effectively leverage model uncertainty.
Discussion. While Gaussian process regression has a computational complexity of $O(N^{3})$ with respect to the number of observations $N$, in our case this complexity is defined w.r.t. to a user’s interaction history, not the entire set of items in the domain. In practice, given the typical sparseness of user interaction data relative to the total item count, this cubic complexity is generally manageable. Where needed, an effective way to tackle the computational demands of GPR is to impose a limit on the length of a user’s interaction history used by GPR4DUR, using, say, the most recent interactions, or a representative set of interactions, if the length exceeds the threshold. This strategy not only helps to control the computational complexity, but helps ensure the model is primarily influenced by the most recent or relevant user interactions, thus making GPR more applicable and efficient in real-world recommender systems.
For future research, the incorporation of collaborative Gaussian processes presents a tantalizing prospect. Our current model primarily focuses on personalization, using the other users solely for the tuning of GPR hyperparameters. However, we postulate that harnessing the power of collaborative learning (e.g., Houlsby et al. (2012)) could further enhance the performance and effectiveness of our approach.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1.2 Preliminaries on Gaussian Process Regression
A.2 Dataset Statistics
The statistics of the three datasets, Amazon (He & McAuley, 2016), MovieLens (Harper & Konstan, 2015), and Taobao [^1] https://tianchi.aliyun.com/dataset/649?lang=en-us, is shown in Table 5. We report the number of users, number of items, number of total interactions, and density of the three datasets, respectively.
@@ A.2 Dataset Statistics
Table 5: The statistics of datasets.
A.3 Ablation Analysis on Different Clustering Options
In this section, we provide the ablation analysis on using various clustering algorithms and different number of clusters on the Taobao dataset. Following Shi et al. (2023), we have selected four distinct clustering algorithms for our comparative analysis: (i) Ward (Ward, 1963): This hierarchical clustering technique focuses on minimizing the variance within each cluster, effectively reducing the overall sum of squared distances across all clusters. (ii) K-Means (Jin & Han, 2010): An iterative clustering algorithm that aims to minimize the sum of squared distances from each data point to the centroid of its assigned cluster, thereby reducing in-cluster variance. (iii) Spectral Clustering (Shi & Malik, 2000): This method performs clustering based on the eigenvectors of the normalized Laplacian, which is computed from the affinity matrix, facilitating the division based on the graph’s inherent structure. (iv) BIRCH (Zhang et al., 1996): A hierarchical clustering approach that efficiently processes large datasets by incrementally constructing a Clustering Feature Tree, which groups data points based on their proximity and other features.
A notable deviation from Shi et al. (2023) is in our handling of categories (interests) for training and inference phases; we maintain consistency in categories across both phases, in contrast to their separate categorization. We further refine our analysis by tuning the number of clusters from 10, 20, and 50.
@@ A.3 Ablation Analysis on Different Clustering Options
IR@50 ED@50 TEI@50 Recall@50 nDCG@50
The higher the better normal-$\Uparrow$ The lower the better normal-$\Downarrow$ The higher the better normal-$\Uparrow$ The higher the better normal-$\Uparrow$ The higher the better normal-$\Uparrow$
@@
Table 6: Result comparison on Taobao dataset using different cluster options. The Interest Coverage (IC) metric is omitted from this report as its value is significantly influenced by the chosen number of clusters, rendering it less relevant for comparative purposes. All reported metrics are evaluated at the top-50.
@@
A.4 Latency and Performance Comparison
Figure 6: Latency and performance comparison across models using the MovieLens dataset. Training and inference latencies are measured in millisecond (ms), and the vertical short lines indicate the standard deviations.
We show the efficiency comparison (with the performance comparison) across models in Fig 6. For a fair comparison, we run all experiments on a single NVIDIA A100 GPU with TensorFlow framework (version 1.12) without any further optimization on the computation. We choose the MovieLens dataset and set the batch as 1.
@@ A.4 Latency and Performance Comparison
In examining the efficiency of various recommendation models, including YoutubeDNN, GRU4Rec, MIND, DGCF, ComiRec, CAMI, and GPR4DUR, notable differences in training and inference times highlight the diverse computational demands across these models. Models like YoutubeDNN and GRU4Rec present a more efficient profile, offering lower latency during both training and inference phases, while perform the worst on the retrieval and ranking tasks. The variability in model efficiency, as seen in the standard deviations of training and inference times, indicates a potential fluctuation in latency that could impact real-world deployment. Our proposed model, GPR4DUR, stands out with its superior performance on the retrieval task (e.g., measured by interest relevance) and the ranking task (e.g., measured by recall). This suggests that for applications valuing enhanced retrieval and ranking accuracy, the trade-off of increased latency may be acceptable. Additionally, GPR4DUR exhibits a low standard deviation in its timing metrics. This consistency in processing times is advantageous for online deployment. It offers a predictable performance, which is crucial for developers when balancing between model efficiency and effectiveness.
@@
A.5 Comparison with VAE-based Method
We notice that the VAE-based method is also a potential baseline, such as VAECF (Liang et al., 2018b), which assumes a distribution over user/item representations. We do not select it as a main baseline for comparison due to (i) VAECF is a parametric model while our model is non-parametric, and (ii) VAECF is not designed for multi-interest retrieval, which is the focus of this work.
To still satisfy this curiosity, we present a comparison between VAECF and GPR4DUR on the Amazon dataset for both the retrieval task and ranking task. As shown in Table 7, we observe that GPR4DUR outperforms VAECF on the retrieval task while being slightly worse than VAECF on the ranking task. Considering the gap on the ranking task compared to VAECF is small and the main focus of this work is the retrieval phase, we can still confirm the effectiveness of our model.
@@ A.5 Comparison with VAE-based Method
IC@50 IR@50 ED@50 TEI@50 Recall@50 nDCG@50
The higher the better normal-$\Uparrow$ The higher the better normal-$\Uparrow$ The lower the better normal-$\Downarrow$ The higher the better normal-$\Uparrow$ The higher the better normal-$\Uparrow$ The higher the better normal-$\Uparrow$
VAECF 0.783 0.482 0.546 -0.082 0.911 0.931
@@
Table 7: Result comparison between VAECF and GPR4DUR on the Amazon dataset.

### v5

@@ 7 Conclusion and Discussion
In this paper, we introduced a density-based user representation model, GPR4DUR, marking the first application of Gaussian process regression for user modeling in multi-interest retrieval. This innovative approach inherently captures dynamic user interests, provides uncertainty-awareness, and proves to be more dimension-efficient than traditional point-based methods. We also establish a new evaluation protocol, developing new metrics specifically tailored to multi-interest retrieval tasks, filling a gap in the current evaluation landscape. Offline experiments validate the adaptability and efficiency of GPR4DUR, and demonstrate significant benefits relative to existing state-of-the-art models. Online simulations further highlight GPR4DUR’s ability to drive user interest exploration by recommendation algorithms that can effectively leverage model uncertainty. The limitations and broader impacts of this study are discussed in Appendix A.9 and Appendix A.10, respectively.
@@ A.1.2 Preliminaries on Gaussian Process Regression
A.2 Details for Retrieval List Generation
After obtaining a DUR $g_{u}$ for $u\in\mathcal{U}$ using GPR, we generate the retrieval list using the posterior $g_{u}(\mathbf{v})$ over all unobserved items. The top-N items with the highest values are selected as our retrieval list. We consider two methods for estimating these values.
The first is Thompson sampling (TS), a probabilistic method that selects items based on posterior sampling [15]. For each item $v$ in a set, we generate a sample from the posterior $g_{u}(\mathbf{v})$, and rank the items using their sampled values. A key advantage of TS is its ability to balance the trade-off between exploration and exploitation, improving the diversity of the recommendation list. The sampling and selection process is:
$s_{u,v}$ $\sim g_{u}(\mathbf{v}),\quad\forall v\notin\mathcal{I}_{u},$ (15)
[…]
where $s_{u,v}$ is the value for item $v$ from user $u$’s sampled function, and $\mathcal{R}_{u}$ is the final list of retrieved items for user $u$.
The second method is Upper Confidence Bound (UCB), a deterministic method that selects items based on their estimated rewards and uncertainties [14, 28]. For each item $v$, we compute its upper confidence bound by adding the mean and a confidence interval derived from the variance of the posterior $g_{u}(\mathbf{v})$; items are ranked using these upper bounds. Unlike TS, UCB tends to aggressively promote items with a high degree of posterior uncertainty, giving a different flavor of diversity in the recommendation list. The selection process is:
$b_{u,v}$ $=\bar{g}_{u}(\mathbf{v})+\beta\cdot\sqrt{\text{var}[{g}_{u}(\mathbf{v})]},\quad\forall v\notin\mathcal{I}_{u},$ (17)
[…]
where $b_{u,v}$ is the upper confidence bound for $v$ w.r.t. $u$’s posterior, and $\beta$ is a hyper-parameter that adjusts the exploration-exploitation trade-off. $\mathcal{R}_{u}$ is the final retrieval list.
A.3 Pretraining Strategies
Necessity of Pretraining Item Embeddings with Auxiliary Loss on Categories. Without augmenting the pre-training loss with categorical information as in Eq. 5, the resulting pre-trained embeddings do not respect category consistency. To illustrate this, we compare pre-trained embeddings with and without categorical information (on MovieLens) and show the results in Fig. 6.
Figure 6: The t-SNE plot for pretrained item embeddings on the MovieLens dataset with and without using auxiliary loss on category information.
We observe that, without categorical constraints, the learned embeddings do not align with the categories. There are still overlap between categories (genres) since each movie may belong to multiple genres. Hence, incorporating an auxiliary training loss based on item categories is essential for interest-coverage analysis based on the embeddings.
Sensitivity to Different Pretraining Strategies. In this work, we use pretrained embeddings for our multi-interest user representation. To ensure a fair comparison, we use the same pretrained embeddings in all other models we compared. To test the impact of the embeddings on performance, we add experiments with other pretraining settings. Specifically, we obtain embeddings using matrix factorization (MF) and using YoutubeDNN with varying history lengths. We use the Amazon dataset as an example and measure performance at top-50. The observation is similar on the other two datasets. Other experiment settings are the same as presented in the paper. We use the learned embeddings with both ComiRec (one of the strongest baselines) and our GPR4DUR. The results, shown in Table 5, demonstrate that embeddings from YoutubeDNN outperform those from MF and that longer history leads to better performance. These results highlight the effectiveness and robustness of our proposed method across different pretraining settings.
[…]
Pretraining setting Retrieval Task Ranking Task
IC@50 $\Uparrow$ IR@50 $\Uparrow$ ED@50 $\Downarrow$ TEI@50 $\Uparrow$ Recall@50 $\Uparrow$ nDCG@50 $\Uparrow$
MF ComiRec 0.684 0.417 0.523 -0.124 0.865 0.851
GPR4DUR 0.887 0.527 0.501 -0.052 0.882 0.871
YoutubeDNN ComiRec 0.709 0.460 0.502 -0.071 0.878 0.870
(max user history=20) GPR4DUR 0.893 0.552 0.426 -0.056 0.892 0.913
YoutubeDNN ComiRec 0.710 0.463 0.493 -0.040 0.887 0.900
(max user history=50) GPR4DUR 0.895 0.560 0.423 -0.039 0.908 0.922
YoutubeDNN ComiRec 0.711 0.465 0.490 -0.039 0.890 0.901
(max user history=100) GPR4DUR 0.896 0.562 0.423 -0.037 0.910 0.925
A.4 Dataset Statistics
The statistics of the three datasets, Amazon [35], MovieLens [16], and Taobao [^1] https://tianchi.aliyun.com/dataset/649?lang=en-us, is shown in Table 6. We report the number of users, number of items, number of total interactions, and density of the three datasets, respectively.
Table 6: The statistics of datasets.
@@ A.4 Dataset Statistics
A.5 Hyperparameter Settings and Sensitivity Analysis for GPR Parameter Tuning
In short, we constantly use the radial basis function (RBF) kernel on different datasets due to its superior performance, where $\sigma$ is unique for each dataset.
The kernel function, $k$, is tuned from the cosine similarity kernel and the RBF kernel. For the RBF kernel, the standard deviation $\sigma$ is tuned from $\{1e^{-2},1e^{-1},1,10,100\}$. For each dataset, we report one metric on the retrieval task and one metric on the ranking task, both evaluated at the top-50. The trend on other metrics is similar.
As shown in Table 7, we observe that the RBF kernel outperforms the cosine similarity kernel across various datasets and metrics because it better captures non-linear relationships and offers flexibility by mapping inputs into an infinite-dimensional space. Additionally, the $\sigma$ parameter in the RBF kernel influences model performance—a larger $\sigma$ yields smoother functions for better generalization over distances, while a smaller $\sigma$ detects finer data variations. Dataset-specific optimal $\sigma$ values correlate with data sparsity. MovieLens’s high density demands a smaller $\sigma$ to avoid overfitting due to capturing detailed variations. In contrast, Taobao’s low density necessitates a larger $\sigma$ for greater generalization, with Amazon’s intermediate sparsity requiring a moderate $\sigma$ for optimal performance.
Table 7: Hyperparameter Settings and Sensitivity Analysis for GPR Parameter Tuning.
GPR Parameters Amazon MovieLens Taobao
[…]
Cosine 0.809 0.874 0.940 0.710 0.518 0.870
RBF ($\sigma=1e^{-2}$) 0.822 0.877 0.963 0.727 0.536 0.872
RBF ($\sigma=1e^{-1}$) 0.883 0.892 0.974 0.730 0.587 0.885
RBF ($\sigma=1$) 0.895 0.908 0.957 0.725 0.591 0.887
RBF ($\sigma=10$) 0.887 0.889 0.954 0.723 0.601 0.906
RBF ($\sigma=100$) 0.890 0.890 0.954 0.721 0.592 0.899
A.6 Ablation Analysis on Different Clustering Options
In this section, we provide the ablation analysis on using various clustering algorithms and different number of clusters on the Taobao dataset. Following paper [49], we have selected four distinct clustering algorithms for our comparative analysis: (i) Ward [50]: This hierarchical clustering technique focuses on minimizing the variance within each cluster, effectively reducing the overall sum of squared distances across all clusters. (ii) K-Means [51]: An iterative clustering algorithm that aims to minimize the sum of squared distances from each data point to the centroid of its assigned cluster, thereby reducing in-cluster variance. (iii) Spectral Clustering [52]: This method performs clustering based on the eigenvectors of the normalized Laplacian, which is computed from the affinity matrix, facilitating the division based on the graph’s inherent structure. (iv) BIRCH [53]: A hierarchical clustering approach that efficiently processes large datasets by incrementally constructing a Clustering Feature Tree, which groups data points based on their proximity and other features.
A notable deviation from paper [49] is in our handling of categories (interests) for training and inference phases; we maintain consistency in categories across both phases, in contrast to their separate categorization. We further refine our analysis by tuning the number of clusters from 10, 20, and 50.
[…]
@@ A.6 Ablation Analysis on Different Clustering Options
[…]
@@
A.7 Latency and Performance Comparison
Table 9: Latency and performance comparison across models on the Taobao dataset, which contains 570K items. Training and inference latencies are measured in millisecond (ms). Standard deviation is shown in parentheses. We see that the cost of inference is on par with other methods.
YoutubeDNN GRU4Rec BERT4Rec gSASRec MIND ComiRec CAMI PIMI REMI GPR4DUR (Ours)
Train (std) 153.08 (1.93) 262.28 (1.20) 926.36 (82.39) 492.10 (25.63) 552.37 (15.50) 575.51 (67.32) 602.38 (84.31) 611.36 (73.02) 732.16 (32.97) 932.00 (16.27)
Inference (std) 1.03 (0.76) 1.44 (0.81) 49.28 (21.47) 20.05 (42.07) 18.62 (52.25) 18.89 (54.57) 38.24 (43.18) 40.57 (39.65) 49.27 (56.33) 63.67 (26.38)
IR@50 0.529 0.533 0.557 0.552 0.552 0.561 0.573 0.566 0.540 0.624
Recall@50 0.875 0.873 0.883 0.884 0.872 0.886 0.891 0.886 0.889 0.906
We show the efficiency comparison (with the performance comparison) across models in Table 9. For a fair comparison, we run all experiments on a single NVIDIA A100 GPU with TensorFlow framework (version 1.12) without any further optimization on the computation. We choose the Taobao dataset and set the batch as 1.
@@ A.7 Latency and Performance Comparison
In examining the efficiency of various recommendation models, notable differences in training and inference times highlight the diverse computational demands across these models. Models like YoutubeDNN and GRU4Rec present a more efficient profile, offering lower latency during both training and inference phases, while perform the worst on the retrieval and ranking tasks. The variability in model efficiency, as seen in the standard deviations of training and inference times, indicates a potential fluctuation in latency that could impact real-world deployment. Our proposed model, GPR4DUR, stands out with its superior performance on the retrieval task (e.g., measured by interest-wise relevance, IR@50) and the ranking task (e.g., measured by recall, Recall@50). This suggests that for applications valuing enhanced retrieval and ranking accuracy, the trade-off of increased latency is acceptable. Additionally, GPR4DUR exhibits a low standard deviation in its timing metrics. This consistency in processing times is advantageous for online deployment. It offers a predictable performance, which is crucial for developers when balancing between model efficiency and effectiveness.
@@
A.8 Comparison with Generative Recommendation Methods
We note that some generative recommendation methods also consider user uncertainty, making them potential candidates for comparison with our method. These methods include VAECF [54], which assumes a distribution over user/item representations, and more recently diffusion-based recommendation models, DiffuRec [55] and DREAM [56]. We do not compare them in the main paper due to (i) different design purpose: these methods are not designed for multi-interest retrieval, which is the focus of this work, and (ii) different model family: VAECF is a parametric model while our model is non-parametric, and diffusion-based methods require an entirely different training procedure and sampling strategies.
To still satisfy this curiosity, we present a comparison between GPR4DUR and the generative recommendation methods on the Amazon dataset for both the retrieval task and ranking task. As shown in Table 10, the results still confirm the effectiveness of our model.
[…]
@@ A.8 Comparison with Generative Recommendation Methods
IC@50 $\Uparrow$ IR@50 $\Uparrow$ ED@50 $\Downarrow$ TEI@50 $\Uparrow$ Recall@50 $\Uparrow$ nDCG@50 $\Uparrow$
VAECF 0.783 0.482 0.546 -0.082 0.909 0.913
DiffuRec 0.834 0.540 0.442 -0.057 0.892 0.916
DREAM 0.812 0.521 0.460 -0.091 0.889 0.917
@@
A.9 Limitations and Future Work
In Sec. 4.5, we discussed methods to reduce the time complexity of traditional Gaussian Process Regression (GPR). Despite these efforts, the training and inference times presented in Table 9 remain slightly higher than those of recent state-of-the-art methods. Further improving the efficiency of our approach would be an intriguing area for future research.
Moreover, the integration of collaborative Gaussian processes offers a promising avenue. Currently, our model focuses primarily on personalization, using other users only for tuning GPR hyperparameters. We hypothesize that leveraging collaborative learning techniques, such as those described in [57], could significantly enhance the performance and effectiveness of our method.
A.10 Broader Impacts
The application of Gaussian Process Regression (GPR) for user modeling in recommendation systems has both positive and negative societal impacts. Below, we outline these impacts in detail.
Positive Societal Impacts. Our approach, GPR4DUR, aims to significantly enhance the personalization and efficiency of recommender systems. By effectively capturing the diverse and dynamic interests of users, GPR4DUR can improve user satisfaction and engagement across various online platforms. This could lead to more tailored content delivery, helping users find relevant and interesting content more quickly and reducing information overload. Furthermore, the uncertainty-aware nature of our method allows for better exploration-exploitation balance, potentially uncovering niche interests and underrepresented content that users might find valuable. This can promote diversity and inclusivity in content consumption, providing a wider range of options to users.
Negative Societal Impacts. Despite these positive aspects, there are potential negative societal impacts associated with the deployment of our method. One major concern is the potential for reinforcing existing biases in recommendation systems. If the training data reflects societal biases, GPR4DUR might inadvertently perpetuate these biases, leading to unfair treatment of certain user groups. Additionally, the improved personalization capabilities might increase the risk of creating filter bubbles, where users are only exposed to content that reinforces their existing preferences, limiting their exposure to diverse viewpoints.
Another potential issue is privacy. As GPR4DUR leverages detailed user interaction histories to model preferences, there is a risk of sensitive information being inferred or misused. Ensuring robust data protection and adhering to privacy standards is crucial to mitigate this risk.
Mitigation Strategies. To address these potential negative impacts, several mitigation strategies can be employed. Firstly, implementing fairness-aware algorithms that explicitly account for and mitigate biases during model training can help ensure equitable recommendations across different user groups. Secondly, incorporating mechanisms to introduce serendipity in recommendations can counteract filter bubble effects, exposing users to a broader range of content. Lastly, adhering to strict data privacy regulations and employing advanced anonymization techniques can safeguard user data and privacy.
Overall, while GPR4DUR has the potential to significantly improve user modeling and personalization in recommender systems, careful consideration and implementation of mitigation strategies are essential to address its broader societal impacts.